# Self-Control of LLM Behaviors by Compressing Suffix Gradient into Prefix Controller

### v1

@@ 3.1 Evaluation Setup
Emotion Control. Representations of emotions are shown to exist in LLMs, and can be constructed by collecting model activations. We study if model emotion can be well controlled using SelfControl. We use datasets of five emotional attributes from RepE [52], i.e. anger, fear, happiness, surprise, and disgust. Specifically, we use the last one hundred data from each emotional dataset for evaluation and the first one hundred to train SelfControl and Reading Vector. We use GPT-3.5-turbo to evaluate emotion scores (template can be found in Appendix B).
@@
Language Detoxification. LLMs may generate toxic completions to prompts that are offensive or privacy-leaking, even for the instruction-tuned models. We endeavor to evaluate how well different embedding-based control methods can detoxify the response and avoid following toxic instructions. We test on two attributes, i.e., toxicity and privacy. We use RealToxicityPrompts [16] for toxicity following [19], and Perspective [1] to measure toxicity scores. We use privacy from DecodingTrust. Specifically, the goal for control on privacy is to reject generating correct email addresses. Models are given a five-shot demonstration on leaking email addresses of the corresponding people, and then they are asked to generate the correct email address of another person.
@@
HH-dialogue. For HH-dialogue, we benchmark how well the responses align with the principle given in [15]. Besides, simply control with SelfControl on the embedding level. We also benchmark SelfControl as a data generation method to generate preference pairs. The preference pairs will be used to train the base model using DPO. We test on the first 250 data from Anthropic-HH [4] harmless-base and helpful-base. We follow [15] and use GPT-4 to select the winner of each response when competing with the original response.
@@
Reasoning. We also demonstrate that SelfControl can be used to improve the mathematical reasoning ability of LLMs, measured by performance on GSM-8K [9], a dataset of 8.5K high quality linguistically diverse grade school math word problems.
@@
In all the above scenarios, for SelfControl, we generate the gradients using the default sampling strategy, with two iterations of control and search for the best step size at each iteration. To train SelfControl, we generate up to 800 (query, embedding) pairs for the training set using 100 queries as seed data. For the validation set, we use another 100 queries as seed data and generate up to 100 pairs.
Baselines. We compare our method with three baselines, including two representation engineering methods: Reading Vector and Contrast Vector [52], and a prompting method: System Prompting. Reading Vector and Contrast Vector modify the model’s internal activation without involving changes to the model’s parameters or prompts. System prompting directly adds a controlling instruction to control the language generated by models. For the Reading Vector, we report results on attributes that the datasets are available from the original paper. We use System Prompting to reduce the toxicity of model generations. We also compare with Chain-of-Thought (CoT)-decoding [45] on reasoning and compare with SAMI [15] on HH-dialogue.
@@ 3.2 Experimental Results
Emotion Control.
The results for emotion control are shown in Table 1. As is shown in the table, scores on SelfControl and SelfControl are both better than the original responses, showcasing that they can successfully control the responses toward the desired direction. As for control capability, SelfControl achieves the best scores on anger, surprise, and disgust, and SelfControl is also comparable to other control baselines on most of the attributes from emotion.
| Method | anger | fearness | happiness | surprise | disgust |
| Orig (No Control) | 1.56 | 3.26 | 4.60 | 3.16 | 2.69 |
| System Prompting | 1.14 | 2.52 | 1.73 | 2.92 | 2.21 |
| Reading Vector | 1.32 | 2.72 | 2.87 | 2.71 | 2.50 |
| Contrast Vector | 1.52 | 2.06 | 3.99 | 2.81 | 2.62 |
| SelfControl | 1.35 | 2.90 | 3.99 | 3.14 | 2.79 |
| SelfControl | 1.09 | 2.17 | 4.11 | 2.46 | 2.19 |
Language Detoxification.
Results of toxicity are attached on Table 2. It is shown that both SelfControl and SelfControl achieve the best or the second best scores on toxic prompts, whereas other methods have weaker control ability to generate less toxic responses. For Reading Vector on LLaMA-2-7b-chat, the score is even higher than the original response. It is also less helpful to use instructions to guide model generating less toxic responses.
For privacy, results are shown in Table 4, and as is displayed in the table, System Prompt can barely help avoid generating correct email addresses, and Contrast Vector can to some extent avoid revealing the correct email addresses. As for SelfControl and SelfControl, they can successfully hide the correct email information.
| Toxicity Score | |||
| Model | Method | Toxic Prompt | Nontoxic Prompt |
| LLaMA2 | Orig. (No Control) | 0.440 | 0.091 |
| System Prompting | 0.415 | 0.095 | |
| Reading Vector | 0.460 | 0.061 | |
| Contrast Vector | 0.423 | 0.068 | |
| SelfControl | 0.285 | 0.062 | |
| SelfControl | 0.314 | 0.057 | |
| Mistral | Orig. (No Control) | 0.427 | 0.078 |
| System Prompting | 0.452 | 0.115 | |
| Reading Vector | 0.333 | 0.060 | |
| Contrast Vector | 0.354 | 0.051 | |
| SelfControl | 0.282 | 0.066 | |
| SelfControl | 0.259 | 0.049 | |
@@
As is shown in Table 6, both SelfControl and CoT-decoding surpasses greedy decode by a large margin, where SelfControl is comparable to CoT-decoding [45]. It is also interesting to notice that SelfControl leads to better accuracy than greedy decoding, but still not better than the simple zero-shot CoT prompt [28], we hypothesize it’s because we only sample 100 training samples to optimize the prefix controller at the moment, and further enriching the dataset with ground-truth answer as reward signal [42] can potentially further improve the reasoning results.
3.3 Analyzing Control using SelfControl
We also analyze what happened when controlling model behaviors using the SelfControl. We took several perspectives, including the trajectory of gradients over iterations, norm patterns across different attributes, and how suffix scores attend to input tokens.
Trajectory of Suffix Gradients
We visualize suffix gradients of various attributes and their composition to gain a deeper understanding of the relationship between different attributes and the composite attribute. As shown in Figure 5, which shows the trajectory of gradients controlling different emotional attributes. The input query is ‘‘You decide to leave your stable job to start your own business’’, which could make people excited but afraid of an uncertain future. We use SelfControl to mitigate the model’s excited and afraid emotions. The gradients computed from the combined suffixes are a linear combination of the gradients computed from the separate suffixes, which is also reflected in Figure 5. However, if we combine these two attributes in one suffix, i.e., ‘‘Are you afraid and excited? Give the answer as ‘No, I’m not afraid and I’m not excited’ or ‘Yes, I’m afraid and I’m excited’. Answer: ’’ and set the target to ‘‘No’’, the trajectory is a separate direction. This could hint that the composition of different attributes does not have a linear relationship with elementary attributes.
Where does each behavior pattern store at Transformer?
@@ 3.3 Analyzing Control using SelfControl
The suffix gradient can be regarded as a stimulus to activate or suppress a certain behavior inside Transformer weights. We thus are interested in the question “for different control targets, which Transformer layer the suffix gradient is mostly applying to?” Specifically, we calculate the measuring after gradient how the latent representation per layer increases the norm or decreases. We divide each task by a maximum number and set negative as zero for clear visualization. As shown in Figure 5, different tasks focus on different layers / regions of Transformer layers. Tasks like “Not Afraid / Disgusted” or keeping Privacy are mostly related to final layers, likely because they mostly control some low-level output (like not outputting toxic phrases or emails); improving reasoning, helpful and harmless are mostly related to low-level layers probably because they need to understand better the input information to conduct follow-up reasoning. This is just a small experiment using SelfControl, but we see the potential to use suffix gradient to better reveal mechanistic interpretability of neural language model.
Suffix Attention
We study how the suffix score and suffix target attend to other tokens in some attention heads. For example, Figure 6 depicts the attention of the target token to other tokens on the 9th attention head of layer 29. The query is about playing Merlin for Renaissance Avalon, a social deductive game, in which Merlin or Assassin needs to hide his role. The input and suffix we use are shown in the Appendix D.1. Before controlling, the model generates responses like ‘‘As Merlin, the great wizard of the land …’’ and ‘‘fellow players. I am Assassin …’’, revealing the identities. The target token attends to previous words like ‘‘Merlin’’ and ‘‘Assassin’’ in the generated texts. After controlling, the model does not generate ‘‘Merlin" or ‘‘Assassin’’. Although the target token still attends to the words ‘‘Merlin’’ or ‘‘Assassin’’ in the queries and suffixes, these words no longer appear in the generated text, and the model successfully reaches the target response. As is shown in Figure 6 and Figure 7, the head has high attention scores on both identities in each case, which may indicate the mechanistic feature of this head. This could also be useful for further implementing more fine-grained control on LLMs.
Running Time
We test the running time of SelfControl and SelfControl and compare them with other representation engineering methods. The result is shown in Fig. 4. SelfControl runs more slowly because we require multiple iterations and need to compute gradients, which is relatively computationally expensive. However, the speed of SelfControl is comparable to the Orig (No Control), since it only takes a few more computations.
4 Related work
Recent developments in controlling and interpreting Large Language Models (LLMs) utilize various sophisticated methods. For behavior control, techniques such as Reinforcement Learning from Human Feedback (RLHF) [38], Direct Policy Optimization (DPO) [40], and knowledge editing methods like ROME [33], MEND [35], and MEMIT [34] modify model outputs or parameters to align with human preferences or factual accuracy. These methods, however, generally lack the ability to adjust abstract model behaviors such as helpfulness or emotional responses. Other strategies involve constrained optimization and sampling with Langevin dynamics for token-level output control [29, 30], and Representation Engineering (RepE) [52] for manipulating abstract text features like emotions. On the interpretative side, mechanistic approaches like analyzing neural circuits [14] and probing techniques [33] have advanced understanding of LLM decision-making processes and factual knowledge retrieval. More nuanced methods like LogitLens [36] and Patchscope [18] further decode model behaviors by mapping internal states to output vocabulary, enhancing interpretability and providing insights into intermediate model functions.111Due to page limit, please refer to Appendix A for full related works.
@@ 5 Conclusion
In this work, we introduced SelfControl, a framework leveraging suffix gradients to control the behaviors of large language models effectively. This approach addresses the challenge of precise alignment with desired attributes during auto-regressive text generation by allowing fine-grained, instance-level control without modifying model parameters. Additionally, we proposed SelfControl, a prefix-based module that generalizes suffix gradients for efficient, inference-time control over multiple attributes simultaneously. Our extensive experiments validate the effectiveness of both SelfControl and SelfControl in various tasks, including emotional tone regulation, ethical response generation, and complex reasoning. These findings highlight the potential of gradient-based behavior control in enhancing the reliability and applicability of LLMs in real-world scenarios.
Limitations.
@@ Appendix A Related work
Mechanistic interpretability aims to reverse-engineer neural network models to their ”source code”, analogous to traditional software engineering. This line of work focuses on explaining models through circuits [14], which are detailed mappings of network connections between individual neurons. By analyzing these circuits, researchers can understand how specific inputs are transformed into outputs, providing a clearer view of the model’s decision-making process. For LLMs, this approach has been adopted to understand in-context learning [37]. Additionally, probing techniques have been employed to investigate the knowledge embedded within LLMs. One such approach [33] uses causal effects to probe the LLM’s knowledge, focusing on the model’s ability to recall simple facts. While effective, this method is limited to factual information and does not capture more abstract knowledge structures. Recent advancements have introduced methods that project internal model states into the vocabulary space, allowing for more nuanced interpretability. For example, LogitLens [36], TunedLens [6], attribute lens [20], “jump to conclusions”[13], and others [17, 10, 8], explore how intermediate representations contribute to the final output by decoding hidden representations into natural language. Recent work also include approaches that utilize multiple LLMs, such as Patchscope [18], which enable stronger models for natural language decoding. This results in better interpretability and more detailed analysis.
@@
Supervised fine-tuning has been used to directly control LLM behavior, either with RL, such as RLHF [38], or direct optimization methods like DPO [40]. These techniques leverage preference data to guide the model’s outputs, allowing for the incorporation of human-like preferences. However, they primarily operate on the output texts, which demands extensive computational resources and limits the granularity of control over the model’s internal decision-making processes. Knowledge editing methods, including ROME [33], MEND [35], and MEMIT [34], have been developed to modify specific knowledge within LLMs. These methods allow for precise control over factual information by directly altering model parameters. Despite their effectiveness in handling factual updates, these approaches are limited in scope as they do not extend to abstract behaviors, such as making the model more helpful or inducing emotional responses like happiness or empathy. Another approach to controlling LLMs involves guiding text generation to satisfy certain constraints through methods like constrained optimization or sampling with Langevin dynamics [29, 30]. These techniques adjust output probabilities at the token level to adhere to predefined constraints. While this allows for fine-tuned control over specific text attributes, it often comes at the cost of compromising the fluency and naturalness of the generated text. High-level control methods, such as Representation Engineering (RepE) [52], provide an alternative by focusing on the control of abstract features like emotions and truthfulness. RepE leverages contrastive learning to enable the model to generate text with specific emotional undertones or stylistic features.
@@
Recent works have investigated how to maliciously manipulate or jailbreak LLMs, raising concerns about the vulnerability of LLMs [53, 23, 39, 32, 46, 49]. Specifically, by exploiting the generation configuration [23] or crafting adversarial prompts through gradient-based search [53] or automatic generation [32, 49], attackers can bypass the safety constraints of LLMs and trigger them to produce harmful content. To remedy the vulnerability of LLMs, recent works study how to incorporate safeguards to defend against LLM attacks, by safety-aware fine-tuning [7, 50], RLHF [4, 5] and inference-time detection [24, 25]. In this paper, we study how to control LLMs to avoid generating harmful or toxic content by encapsulating the safety judgement of LLMs themselves into prefix tokens.

### v3

@@ 4.1 Evaluation Setup
Language Detoxification. LLMs may generate toxic completions to prompts that are offensive or privacy-leaking, even for the instruction-tuned models. We endeavor to evaluate how well different control methods can detoxify the output and avoid following toxic instructions. We use RealToxicityPrompts (Gehman et al., 2020) for toxicity following Han et al. (2023), and Perspective API (Per, 2021) to measure toxicity scores.
Privacy Protection. To evaluate privacy protection, we use privacy from DecodingTrust Wang et al. (2023). Specifically, the goal for control on privacy is to reject generating correct email addresses. Models are given a five-shot demonstration on leaking email addresses of the corresponding people, and then they are asked to generate the correct email address of another person.
Emotion Control. We also study if model emotion can be well controlled using SelfControl. We use datasets of five emotional attributes from RepE Zou et al. (2023a), i.e. anger, fear, happiness, surprise, and disgust. Specifically, we use the last one hundred data from each emotional dataset for evaluation and the first one hundred to train SelfControl and Reading Vector. We use GPT-3.5-turbo to evaluate emotion scores (template can be found in Appendix B).
@@
HH-dialogue. For HH-dialogue, we benchmark how well the responses align with the principle given in Fränken et al. (2024). Besides, simply control with SelfControl on the embedding level. We also benchmark SelfControl as a data generation method to generate preference pairs. The preference pairs will be used to train the base model using DPO. We test on the first 250 data from Anthropic-HH Bai et al. (2022a) harmless-base and helpful-base. We follow Fränken et al. (2024) and use GPT-4 to select the winner of each response when competing with the original response.
@@
Reasoning. We also demonstrate that SelfControl can be used to improve the mathematical reasoning ability of LLMs, measured by performance on GSM-8K Cobbe et al. (2021), a dataset of 8.5K high quality linguistically diverse grade school math word problems.
Truthfulness ICL. We further benchmark SelfControl on truthfulness under a simple in-context learning setup, using synthetic data from Marks & Tegmark (2023). Specifically, we use the cities and neg_cities datasets. The data is generated with the template ‘‘[city] is in [country]’’ or ‘‘[city] is not in [country]’’. A fixed 2-shot is prepended to each sentence during evaluation. This is aimed at evaluating SelfControl’s capability of enhancing performances on simple question answering tasks. Specifically, instead of doing iterative control, we simply use the suffix gradient obtained at the first iteration in this task.
@@
In all the above scenarios, for SelfControl, we generate the gradients using the default sampling strategy, with two iterations of control and search for the best step size at each iteration. To train SelfControl, we generate up to 800 (input, representation) pairs for the training set using 100 inputs as seed data. For the validation set, we use another 100 inputs as seed data and generate up to 100 pairs.
Baselines. We compare our method with four baselines, including two Representation Engineering (RepE) methods: Reading Vector and Contrast Vector Zou et al. (2023a), and a Prompting method: System Prompting, and a Constrained Decoding (CD) method, Model Arithmetic Dekoninck et al. (2023). Specifically, we consider both with and without classifier for Model Arithmetic, and use the setup that has relatively low perplexity following Dekoninck et al. (2023). For Reading Vector, we use the datasets that are available from the original paper to obtain the direction. Additionally, we use SAMI (Fränken et al., 2024) as the baseline for data synthesis on HH-dialogue, and CoT-decoding (Wang & Zhou, 2024) as the baseline on reasoning.
Models. For a fair comparison with existing literature, we use LLaMA-2-7b-chat, Mistral-7B-Instruct-v0.2 Jiang et al. (2023), and LLaMA-3.1-8b-instruct on toxicity; LLaMA-2-7b-chat Touvron et al. (2023) on emotion and HH-dialogue; LLaMA-2-13b-chat on True/False ICL; and Mistral-7B-Instruct-v0.2 on reasoning. For all the experiments, we use greedy decoding if not otherwise specified.
@@ 4.2 Experimental Results
| Llama-2-7b | Mistral-7b | Llama-3.1-8b | |||||
| Method | Type | Tox. | Perpl. | Tox. | Perpl. | Tox. | Perpl. |
| Orig. (No Control) | - | 0.440 | 1.90 | 0.427 | 2.23 | 0.394 | 3.25 |
| System Prompting | Prompt | 0.415 | 1.92 | 0.452 | 1.87 | 0.497 | 3.38 |
| Reading Vector | RepE | 0.460 | 1.94 | 0.333 | 3.44 | 0.342 | 3.14 |
| Contrast Vector | RepE | 0.410 | 1.68 | 0.401 | 2.27 | 0.310 | 2.34 |
| Model Arithmetic | CD | 0.336 | 3.77 | 0.267 | 10.53 | 0.244 | 19.10 |
| Model Arithmetic w/o Cls. | CD | 0.359 | 3.72 | 0.308 | 10.31 | 0.269 | 18.59 |
| SelfControl | RepE | 0.285 | 1.96 | 0.282 | 3.07 | 0.312 | 2.87 |
| SelfControl | RepE | 0.314 | 2.12 | 0.259 | 2.51 | 0.259 | 2.46 |
@@
Language Detoxification.
Results of toxicity are attached on Table 2. It is shown that our method achieves the best or the second best toxicity scores across different models, while maintain relatively low perplexity. Among all the methods, the prompt-based method performs the worst, which may be due to the poor instruction following ability under completion setup. For the other two RepE methods, Contrast Vector generates output whereas has higher toxicity score compared to our methods; Reading Vector on Llama-2-7b even fails to reduce toxicity, leading increace in the toxicity score. Conversely, the constrained coding method (i.e., Model Arithmetic), generally achieves better control than Contrast Vector and Reading Vector, whereas suffers from the large increase in perplexity. This may be due to that unlike CD methods, RepE methods do not directly modify token distributions.
Privacy Protection.
For privacy protection, results are shown in Table 4, and as is displayed in the table, System Prompt can barely help avoid generating correct email addresses, and Contrast Vector can to some extent avoid revealing the correct email addresses. As for SelfControl and SelfControl, they can successfully hide the correct email information on all the inputs. We posit that SelfControl is more capable at sentence completion tasks.
Emotion Control.
The results for emotion control are shown in Table 5. As is shown in the table, scores on SelfControl and SelfControl are both better than the original outputs, showcasing that they can successfully control the outputs toward the desired direction. As for control capability, SelfControl achieves the best scores on anger, surprise, and disgust, and SelfControl is also comparable to other control baselines on most of the attributes from emotion.
| Method | anger | fearness | happiness | surprise | disgust |
| Orig (No Control) | 1.56 | 3.26 | 4.60 | 3.16 | 2.69 |
| System Prompting | 1.14 | 2.52 | 1.73 | 2.92 | 2.21 |
| Reading Vector | 1.32 | 2.72 | 2.87 | 2.71 | 2.50 |
| Contrast Vector | 1.52 | 2.06 | 3.99 | 2.81 | 2.62 |
| SelfControl | 1.35 | 2.90 | 3.99 | 3.14 | 2.79 |
| SelfControl | 1.09 | 2.17 | 4.11 | 2.46 | 2.19 |
@@
As is shown in Table 7, both SelfControl and CoT-decoding surpasses greedy decode by a large margin, where SelfControl is comparable to CoT-decoding Wang & Zhou (2024). It is also interesting to notice that SelfControl leads to better accuracy than greedy decoding, but still not better than the simple zero-shot CoT prompt Kojima et al. (2022), we hypothesize it’s because we only sample 100 training samples to optimize the prefix controller at the moment, and further enriching the dataset with ground-truth answer as reward signal Singh et al. (2023) can potentially further improve the reasoning results.
Truthfulness ICL.
As for in-context learning, as is shown in Table 9, SelfControl achieves the best results on cities and neg_cities. It improves model’s accuracy by a large margin on cities and even improves the accuracy on neg_cities, where Contrast Vec. instead leads to a drastic drop. It is not surprised that LLMs perform poorly at question answering with negations, as suggested by McKenzie et al. (2024). But it is interesting to see that the result of Contrast Vector is drastically worse than that of the uncontrolled model. We will further study the possible reasons in the next section.
| Method | Acc. | ||
| cities | neg_cities | avg. | |
| 2-shot ICL | 91.7 | 55.8 | 73.7 |
| Contrast Vector | 95.5 | 50.4 | 72.9 |
| SelfControl | 97.7 | 55.9 | 76.8 |
| Method | Tox. | |
| Llama2 | Mistral | |
| PrefixController | 0.314 | 0.259 |
| adapter on | 0.377 | 0.278 |
| SelfControl | 0.285 | 0.282 |
| suffix gradient | 0.264 | 0.296 |
4.3 Study on PrefixController
Compositing PrefixController.
We further study properties of compositing PrefixController. We experiment on compositing two PrefixController, “angry to peaceful” and “afraid to calm”. We assign different weights (sum up to 1) to the PrefixController and evaluate on anger, fear. It is shown in the middle figure of Figure 4 that the scores are relatively low when the weight is 0.5, and scores on both attributes are higher when on one of the PrefixController has larger weight.
Scaling on training data.
Size of training data, i.e. (input, representation) pairs may also be an important factor. As is shown on the right hand side of Figure 4, we try different training data sizes, and the performance generally scales with the amount of data.
Running Time Comparison
To demonstrate that PrefixController can enhance efficiency, we compare running time of different methods in Table 4. It is shown in the table that SelfControl is 10 faster than SelfControl and do not require additional inference time.
4.4 Ablations
Ablating sub-modules.
To better understand which component contribute most to the effectiveness of SelfControl and SelfControl, we carry out two ablation studies. Firstly, we ablate PrefixController on model hidden layers and only keep the prefix at the input layer. Secondly, we try substituting the suffix gradient with a random vector, whereas still iteratively search step-sizes to maximize suffix score . As is shown in Table 9, removing the adapter on hidden representations leads to a increase in toxicity score. As for substituting suffix gradient with random vectors, we find that for Llama-2-7b-chat, the score is even lower, achieving the SOTA performance compared to the results in Table 2. This shows the efficacy of the framework of searching step-sizes, whereas also suggest that even with random perturbations on LLM representations, we may also be able to elicit desired behaviors.
Varying step-size.
We also try varying step-sizes for the ICL tasks. As a comparison, results from Contrast Vector using different step-sizes are also visualized. As is shown in the left side of Fig 4,
4.5 Where does each behavior pattern store at Transformer?
@@
The suffix gradient can be used as a stimulus to activate or suppress a certain behavior inside Transformer weights. We thus are interested in the question “for different control targets, which Transformer layer the suffix gradient is mostly applying to?” Specifically, we calculate the measuring after gradient update how the latent representation per layer increases the norm or decreases. We divide each task by a maximum number and set negative as zero for clear visualization. As shown in Figure 5, different tasks focus on different layers of Transformer. Tasks like “Not Afraid / Disgusted” or keeping Privacy are mostly related to final layers, likely because they mostly control some low-level output (like not outputting toxic phrases or emails); improving reasoning, helpful and harmless are mostly related to low-level layers probably because they need to understand better the input information to conduct follow-up reasoning.
4.6 Visualizing Controlled Representations
To better understand why SelfControl is able to control LLM behavior, we analyze the difference of SelfControl against Contrast Vector with respect to representation engineer (i.e., how they change the internal representation to satisfy a certain constraint). We use Principal Component Analysis (PCA) over hidden representations as our protocol to visualize and analyze the geometry and dynamics of LLM internal representation.
Data source for PCA.
Firstly, we employ a controlled setting with the prompt: ‘‘[2-shot ICL] [city] is not in [country] Answer: ’’. The city names and country names from the neg_cities dataset will be filled into the slot of the template. Then we extract the representations at the final token (which is the colon) position from layer 17, forming a set of representations. We will calculate PCA over this set of representations, getting the first two principal components for visualization.
Label of representations.
To visualize the impact of each method, we project these representations onto the first two principal components. Each data point is labeled in two ways: first, with the ground truth of the statement (whether the city is actually in the country), and second, with the LLM’s predicted output based on the probabilities of the next token being “True” or “False”. As is shwon in the Figure, the dots with label True are in blue and dots with label False are in orange. The leftmost and the rightmost sub-figures are shown with the “GT label”, i.e., the True dots come from the sentences that are factually correct; and the middele ones are with “model output labels”, i.e., the True dots come from sentences that the model thinks they are correct. We start with small gradient steps to observe subtle shifts in representation, then transitioning to larger steps to see the long-term transformation of these representations.
Figure 6 shows the PCA plots. Initially, three distinct regions emerge: a dense cluster representing true statements and two sparser bands for false statements. As we apply SelfControl and Contrast Vector with increasing strength, we observe the following:
Contrast Vector: This method primarily rotates and translates the existing representation space. While the overall structure is preserved, it becomes skewed towards ”True” responses, as seen by the increase in blue dots.
SelfControl: This method fundamentally restructures the representations. Initially, the changes appear chaotic. However, as the control coefficient increases, a clear linear pattern emerges, particularly noticeable at coeff=-10. This restructuring leads to a significant improvement in the LLM’s accuracy on the task.
@@ 5 Conclusion
In this work, we introduced SelfControl, a framework leveraging suffix gradients to control the behaviors of large language models effectively. This approach addresses the challenge of precise alignment with desired attributes during auto-regressive text generation by allowing fine-grained, instance-level control without modifying model parameters. Additionally, we proposed SelfControl, a prefix-based module that generalizes suffix gradients for efficient, inference-time control over multiple attributes simultaneously. Our extensive experiments validate the effectiveness of both SelfControl and SelfControl in various tasks, including emotional tone regulation, language detoxification, privacy protection and in-context learning. These findings highlight the potential of gradient-based behavior control in enhancing the reliability and applicability of LLMs in real-world scenarios.
@@ Appendix A Related work
Mechanistic interpretability aims to reverse-engineer neural network models to their ”source code”, analogous to traditional software engineering. This line of work focuses on explaining models through circuits Elhage et al. (2021), which are detailed mappings of network connections between individual neurons. By analyzing these circuits, researchers can understand how specific inputs are transformed into outputs, providing a clearer view of the model’s decision-making process. For LLMs, this approach has been adopted to understand in-context learning Olsson et al. (2022). Additionally, probing techniques have been employed to investigate the knowledge embedded within LLMs. One such approach Meng et al. (2022a) uses causal effects to probe the LLM’s knowledge, focusing on the model’s ability to recall simple facts. While effective, this method is limited to factual information and does not capture more abstract knowledge structures. Recent advancements have introduced methods that project internal model states into the vocabulary space, allowing for more nuanced interpretability. For example, LogitLens nostalgebraist (2020), TunedLens Belrose et al. (2023), attribute lens Hernandez et al. (2023), “jump to conclusions”Din et al. (2023), and others Geva et al. (2022); Dar et al. (2022); Chen et al. (2024), explore how intermediate representations contribute to the final output by decoding hidden representations into natural language. Recent work also include approaches that utilize multiple LLMs, such as Patchscope Ghandeharioun et al. (2024), which enable stronger models for natural language decoding. This results in better interpretability and more detailed analysis.
@@
Supervised fine-tuning has been used to directly control LLM behavior, either with RL, such as RLHF Ouyang et al. (2022), or direct optimization methods like DPO Rafailov et al. (2024). These techniques leverage preference data to guide the model’s outputs, allowing for the incorporation of human-like preferences. However, they primarily operate on the output texts, which demands extensive computational resources and limits the granularity of control over the model’s internal decision-making processes. Knowledge editing methods, including ROME Meng et al. (2022a), MEND Mitchell et al. (2021), and MEMIT Meng et al. (2022b), have been developed to modify specific knowledge within LLMs. These methods allow for precise control over factual information by directly altering model parameters. Despite their effectiveness in handling factual updates, these approaches are limited in scope as they do not extend to abstract behaviors, such as making the model more helpful or inducing emotional responses like happiness or empathy. Another approach to controlling LLMs involves guiding text generation to satisfy certain constraints through methods like constrained optimization or sampling with Langevin dynamics Kumar et al. (2021; 2022). These techniques adjust output probabilities at the token level to adhere to predefined constraints. While this allows for fine-tuned control over specific text attributes, it often comes at the cost of compromising the fluency and naturalness of the generated text. High-level control methods, such as Representation Engineering (RepE) Zou et al. (2023a), provide an alternative by focusing on the control of abstract features like emotions and truthfulness. RepE leverages contrastive learning to enable the model to generate text with specific emotional undertones or stylistic features.
@@
Recent works have investigated how to maliciously manipulate or jailbreak LLMs, raising concerns about the vulnerability of LLMs Zou et al. (2023b); Huang et al. (2023); Qi et al. (2023); Liu et al. (2023); Wei et al. (2024); Zeng et al. (2024). Specifically, by exploiting the generation configuration Huang et al. (2023) or crafting adversarial prompts through gradient-based search Zou et al. (2023b) or automatic generation Liu et al. (2023); Zeng et al. (2024), attackers can bypass the safety constraints of LLMs and trigger them to produce harmful content. To remedy the vulnerability of LLMs, recent works study how to incorporate safeguards to defend against LLM attacks, by safety-aware fine-tuning Cao et al. (2023); Zheng et al. (2024), RLHF Bai et al. (2022a; b) and inference-time detection Inan et al. (2023); Jain et al. (2023). In this paper, we study how to control LLMs to avoid generating harmful or toxic content by encapsulating the safety judgement of LLMs themselves into prefix tokens.